# Contrastive Chain-of-Thought Prompting

### v1

@@ Abstract
Contrastive Chain-of-Thought Prompting
Grant Kruttschnitt           Jay Shim           Alyssa Ma           Daniel Kim Benjamin Chek           Athul Anand                     Kevin Zhu           Sean O’Brien Algoverse AI Research sean@algoverse.us, kevin@algoverseacademy.com
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing by leveraging vast amounts of data to generate human-like text. These models have gained attention for their ability to perform a wide range of language tasks, from text generation and translation to question answering and summarization. One such technique that has propelled the performance of LLMs is Chain-of-Thought (CoT) prompting (Wei et al., 2022b) . By applying CoT prompting, researchers have found substantial improvement in various tasks such as arithmetic and common sense reasoning, surpassing the efficacy of traditional methods focused on scaling model size (Wang et al., 2023; Wei et al., 2023). In a similar vein, Context-Aware Decoding (CAD) emerges as another approach to augment LLM performance (Liu et al., 2021; Li et al., 2023a; O’Brien and Lewis, 2023), priotizing relevant contextual information over pre-existing knowledge (Shi et al., 2023). While these advancements underscore significant progress, an ongoing question persists regarding the efficacy of combining CAD with CoT or with amateur prompts, which typically contain less context.
@@ 2.1 Background
We take a query x which we feed to our language model $\theta$, to generate a response y using greedy decoding. The response is sampled autoregressively from two probability distributions given by query $x$, CoT prompt c, and an amateur model (3.2.1) with less context amt (Shi et al., 2023). Using a contrastive strength of $\alpha$, we create a new distribution:
@@ 2.2 Context-Aware Decoding
For each method we utilize two different prompts: an 8-shot CoT prompt, and an amateur prompt. To identify the next logit and thus the next token to add autoregressively, we use the contrastive function below proposed by Shi et al. (2023):
@@
The values of $\alpha$ are as follows, $0\leq\alpha\leq 1$, with higher alpha values indicating stronger reliance on prompt c, the CoT prompt. From the adjusted logits, we apply greedy decoding to generate the subsequent character. In our analysis, a single model was used to generate probabilities for both CoT and amateur prompts, but different models may also be used.
@@ 3.2 Models and Prompting
We performed our experiments using Phi-1.5, Mistral 7B, and GPT-3.5 with version 3.5-turbo-0125. In each of our experiments, we contrasted the logits obtained from a 8-shot CoT prompt with the amateur. The exemplars in the few-shot prompts are taken from the corresponding dataset used for evaluation.
The expert prompt shared across all experiments is displayed below.
@@ 3.2.1 Expert vs Amateur 1
In our main method, the amateur is given no context. By doing so, we increase the potential of our context-aware decoding and follow the formulation of (Sanchez et al., 2023)’s method.
The effectiveness on this method is that this will cause it to hopefully rely on the context much more, since the tokens that do rely on it will be much more likelier to occur.
@@ 3.3 Hyperparameter Search
Measuring performance on Mistral-7B (Jiang et al., 2023) using the CommonSenseQA dataset (Talmor et al., 2019), we found an optimal $\alpha$ value of $0.8$ (see  Table 1). This is a constant we used across all datasets and models. In addition, we decided to use 0.5, because although it did not have the second highest accuracy, we wanted another alpha away from 0.8. That would allow us to see a more pronounced difference between the alphas.
@@ 4 Results
GSM8K notably shows a significant decrease in accuracy on Phi-1.5 using CoT contrast. This could be due to contamination in the Phi-1.5B dataset (Zhang et al., 2024)
@@
Table 5 shows an interesting property of the outputs, as it appears the baseline evaluates its expressions with lower accuracy than most of the other experiments, but still results in a higher accuracy. After conducting analysis on model output for each of the experiments, it appears as though, for the experiments, the mathematical calculations may be higher in accuracy in the Amateurs, but the expressions themselves contain incorrect numerical values which may indicate an incorrect translation from the word questions to expressions.
The overall results demonstrate that contrasting with CoT prompting and amateur models is mostly successful when addressing common-sense multiple choice questions.
@@
Table 3: GPT-3.5 Results on Amateur 1 on 0.8 Alpha. It does not perform as well on GPT, with only one of the CCoT having a better result.
@@ 5 Related Work
The first direction that this paper relates to is prompting. The last few years have given way to major improvements in input part of prompt instructions. This is apparent with the popularity of few-shot prompting by (Brown et al., 2020), automatically learning prompts (Lester et al., 2021), or providing model prompts by describing tasks (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022). This idea leads to our paper’s relation to Chain-of-Thought Prompting.
@@ 6 Conclusion
Our findings highlight the impact of integrating expert and amateur models in Context-Aware Decoding across diverse datasets. This study highlights the need for future research prioritizing context-sensitive applications in language modeling. Further, there is a pressing need to explore the scalability of these techniques and their relevance across even broader datasets and across differing models and sizes. Future investigations could delve into testing a wider range of alpha values to unveil potentially more substantial results, aiding our understanding of the nuances within context-driven language modeling.
@@ 7 Limitations
Our analysis could be limited in that the difference in strength between the expert and amateur models is too inconsequential. That means that there is a possibility that the models themselves have generated outputs that are not dissimilar enough to impact results.
Another limitation specifically on the Phi 1.5 GMS8K results could be the fact that Phi 1.5 is possibly contaminated with GSM8K (Zhang et al., 2024). This explains why our trials for Phi - 1.5 produced many unfinished and illogical answers when responding to input prompts, yet performed much better on the baseline compared on Amateur 1 on GSM8K (Table 2).
@@
Among the three, only GPT-3.5 Turbo is an instruction-tuned model, meaning that it’s ability to reason through questions may not be a result of CoT contrasting but rather because it already contains such capabilities. There also might be things similar to our contrasting methods already hidden in their instructions, which would make our methods less effective.

### v2

# Chain-of-Thought Augmentation with Logit Contrast for Enhanced Reasoning in Language Models

@@ Abstract
Chain-of-Thought Augmentation with Logit Contrast for Enhanced Reasoning in Language Models
Jay Shimthanks: Lead Authors with Equal Contribution. Order chosen randomly.           Grant Kruttschnitt11footnotemark: 1           Alyssa Ma           Daniel Kim Benjamin Chek           Athul Anand                     Kevin Zhuthanks: Senior Author.           Sean O’Brien22footnotemark: 2 Algoverse AI Research sean@algoverse.us, kevin@algoverse.us
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing by leveraging vast amounts of data to generate human-like text. These models have gained attention for their ability to perform a wide range of language tasks, from text generation and translation to question answering and summarization. One such technique that has propelled the performance of LLMs is Chain-of-Thought (CoT) prompting (Wei et al., 2022b) . By applying CoT prompting, researchers have found substantial improvement in various tasks such as arithmetic and common sense reasoning, surpassing the efficacy of traditional methods focused on scaling model size (Wang et al., 2023; Wei et al., 2023). In a similar vein, Context-Aware Decoding (CAD) emerges as another approach to augment LLM performance (Liu et al., 2021; Li et al., 2023a; O’Brien and Lewis, 2023), prioritizing relevant contextual information over pre-existing knowledge (Shi et al., 2023). While these advancements underscore significant progress, an ongoing question persists regarding the efficacy of combining CAD with CoT or with amateur prompts, which typically contain less context.
@@ 2.1 Background
We take a query x which we feed to our language model $\theta$, to generate a response y using greedy decoding. The response is sampled autoregressively from two probability distributions given by query $x$, CoT prompt c, and an amateur prompt (3.2.1) with less context amt (Shi et al., 2023). Using a contrastive strength of $\alpha$, we create a new distribution:
@@ 2.2 Context-Aware Decoding
For each method we utilize two different prompts: an 8-shot CoT prompt which we call an expert prompt, and an amateur prompt. To identify the next logit and thus the next token to add autoregressively, we use the contrastive function below proposed by Shi et al. (2023):
@@
The values of $\alpha$ are as follows, $0\leq\alpha\leq 1$, with higher alpha values indicating stronger reliance on prompt c, the CoT prompt. From the adjusted logits, we apply greedy decoding to generate the subsequent token. In our analysis, a single model was used to generate probabilities for both CoT and amateur prompts, but different models may also be used.
@@ 3.2 Models and Prompting
We performed our experiments using Phi-1.5 (Li et al., 2023c), Mistral 7B (Jiang et al., 2023a), and GPT-3.5 (OpenAI et al., 2024) with version 3.5-turbo-0125. The reason we chose these models is that they were varied. The reason we used Phi-1.5 was because it was very low parameters compared to the other two models. We used Mistral, because it had more parameters, so it was more accurate. The reason we used GPT 3.5 was to compare it with the other two open-source models, and because it was more parameters than Mistral. However, our use of GPT 3.5 was limited, because it cost a lot, so we were only able to run baseline and one Amateur on one Alpha. In each of our experiments, we contrasted the logits obtained from a 8-shot CoT prompt with the amateur. The exemplars in the few-shot prompts are taken from the corresponding dataset used for evaluation.
To clarify, for each of these, the amateur prompt is different, but has a limitation which should make it worse. We contrast it against the expert prompt, which had the same format across all experiments, as seen below.
@@ 3.2.1 Expert vs Amateur 1
In our main method, the amateur is given no context, not even the question. By doing so, we increase the potential of our context-aware decoding and follow the formulation of Sanchez et al. (2023)’s method.
The effectiveness on this method is that this will cause it to hopefully rely on the chain of thought process, and the actual question much more, since the tokens that do rely on it will be much more likelier to occur.
@@ 3.3 Hyperparameter Search
Measuring performance on Mistral 7B (Jiang et al., 2023b) using the CommonSenseQA dataset (Talmor et al., 2019), we found an optimal $\alpha$ value of $0.8$ (see  Table 1). This is a constant we used across all datasets and models. In addition, we decided to use 0.5, because although it did not have the second highest accuracy, we wanted another alpha away from 0.8. That would allow us to see a more pronounced difference between the alphas.
@@ 4 Results
The reason that we did not put GPT-3.5 into table 2 was because of prohibitive costs of GPT tokens, we could not run the same experiments on GPT that we did on Mistral and Phi.
In Table 3, GSM8K notably shows a significant decrease in accuracy on Phi-1.5 using CoT contrast. This could be due to contamination in the Phi-1.5B dataset (Zhang et al., 2024)
@@
Table 5 shows an interesting property of the outputs, as it appears the baseline evaluates its expressions with lower accuracy than most of the other experiments, but still results in a higher accuracy. After conducting analyses on model output for each of the experiments, it appears as though, for the experiments, the mathematical calculations may be higher in accuracy in the Amateurs, but the expressions themselves contain incorrect numerical values which may indicate an incorrect translation from the word questions to expressions.
The overall results demonstrate that contrasting with CoT prompting and amateur prompting is mostly successful when addressing common-sense multiple choice questions.
@@
Table 3: GPT-3.5 Results on Amateur 1 on 0.8 Alpha. GPT-3.5 use was limited due to expenses so testing was only possible with one Alpha value. It does not perform as well on GPT, with only one of the CCoT having a better result.
@@ 5 Related Work
The first direction that this paper relates to is prompting. The last few years have given way to major improvements in prompting. This is apparent with the popularity of few-shot prompting by (Brown et al., 2020), automatically learning prompts (Lester et al., 2021), or providing model prompts by describing tasks (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022). This idea leads to our paper’s relation to Chain-of-Thought Prompting.
@@ 6 Conclusion
Our findings highlight the impact of integrating expert and amateur prompts in Context-Aware Decoding across diverse datasets. This study highlights the need for future research prioritizing context-sensitive applications in language modeling. Further, there is a pressing need to explore the scalability of these techniques and their relevance across even broader datasets and across differing models and sizes. Future investigations could delve into testing a wider range of alpha values to unveil potentially more substantial results, aiding our understanding of the nuances within context-driven language modeling.
@@ 7 Limitations
Our analysis could be limited in that the difference in strength between the expert and amateur prompting is too inconsequential. That means that there is a possibility that the models themselves have generated outputs that are not dissimilar enough to impact results.
Another limitation specifically on the Phi-1.5 GMS8K results could be the fact that Phi-1.5 is possibly contaminated with GSM8K (Zhang et al., 2024). This explains why our trials for Phi-1.5 produced many unfinished and illogical answers when responding to input prompts, yet performed much better on the baseline compared on Amateur 1 on GSM8K (Table 2).
@@
Among the three, only GPT-3.5 Turbo is an instruction-tuned model, meaning that it’s ability to reason through questions may not be a result of CoT contrasting but rather because it already contains such capabilities. There also might be factors similar to our contrasting methods already hidden in their instructions, which would make our methods less effective.